The perimeter of a rectangle of a wood design is 90 feet. The design length is 5 feet less than four times it's width. Determine the dimensions of the wood design
Answer by addingup(3677)

2l+2w = 90 (1)
and
l = 4w-5 Use this value for l in equation (1):
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
2(4w-5)+2w = 90
8w-10+2w = 90
10w = 100
w = 10 so this is the width. And the length:
l = 4w-5 = 4(10)-5 = 40-5 = 35 The length is 35
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
2(35)+2(10) = 70+20 = 90 Correct.